RDJPGCOM(1) RDJPGCOM(1) NAME rdjpgcom−displaytextcommentsfromaJPEGfile SYNOPSIS rdjpgcom[−verbose][filename] DESCRIPTION rdjpgcomreadsthenamedJPEG/JFIFfile,orthestandardinputifnofileisnamed,andprints anytextcommentsfoundinthefileonthestandardoutput. TheJPEGstandardallows"comment"(COM)blockstooccurwithinaJPEGfile.Althoughthe standarddoesn’tactuallydefinewhatCOMblocksarefor,theyarewidelyusedtoholdusersuppliedtextstrings.Thisletsyouaddannotations,titles,indexterms,etctoyourJPEGfiles, andlaterretrievethemastext.COMblocksdonotinterferewiththeimagestoredintheJPEG file.ThemaximumsizeofaCOMblockis64K,butyoucanhaveasmanyofthemasyoulikein oneJPEGfile. OPTIONS −verbose CausesrdjpgcomtoalsodisplaytheJPEGimagedimensions. Switchnamesmaybeabbreviated,andarenotcasesensitive. HINTS rdjpgcomdoesnotdependontheIJGJPEGlibrary.ItssourcecodeisintendedasanillustrationoftheminimumamountofcoderequiredtoparseaJPEGfileheadercorrectly. SEEALSO cjpeg(1),djpeg(1),wrjpgcom(1) AUTHOR IndependentJPEGGroup 8July1994 1